Multi-Variable Calculus Assignment

due Friday, March 16, 2001

Exercises

1. Find the largest value of the polynomial t^{3} − 3t + 1 on the interval −2 ≤ t ≤ 2.

2. Find the maximum value of the function

f(x, y) = 2x − 3y

on the disk x^{2} + y^{2} ≤ 4.